# Computing solvation free energies of small molecules with first principles accuracy

### v4

@@ I Introduction
[…]
@@
$U(\lambda,r)=4\epsilon\lambda^{n}\Bigg{[}\Bigg{.}\left(\alpha_{LJ}(1-\lambda)^{m}+\left(\frac{r}{\sigma}\right)^{6}\right)^{-2}-\\ \left(\alpha_{LJ}(1-\lambda)^{m}+\left(\frac{r}{\sigma}\right)^{6}\right)^{-1}\Bigg{.}\Bigg{]}$ (3)
@@
Although MLPs can deliver significant accuracy improvements compared to empirical forcefields, their computational expense and the high sampling requirements of free energy calculations have meant that their application has been mostly restricted to corrective perturbations [32, 33, 34]. In this approach, the intramolecular interactions of a subset of atoms (usually the organic molecule) are modelled by the MLP, while the remaining bonded interactions and all nonbonded interactions are modelled by the empirical forcefield. Although there is some evidence that this approach can modestly improve the accuracy of RBFE calculations, it does not address the parametrisation of nonbonded interactions, which are known to be crucial for obtaining accurate free energies [33]. Indeed, it has recently been shown that, for hydration free energies, MLP corrections with mechanical embedding fail to provide statistically significant accuracy improvements [34]. It is worth noting that more sophisticated electrostatic embedding ML/MM schemes have recently been proposed [35, 36]. However, the performance of these methods in corrective binding free energy calculations has not yet been established.
@@ II Results & Discussion
II.1 Alchemical simulations with machine learned forcefields
@@ II.1.1 Softcore dimer curves
To construct the softcore dimer curves, we first computed dimer curves with DFT at the same $\omega$B97M-D3BJ/def2-TZVPPD level of theory as the SPICE dataset [44] for all combinations of the 10 elements covered by MACE-OFF [25]. For each pair, we started by matching the value and gradient of the DFT force with a polynomial of the form $ax^{10}+b$ that approaches a constant value as $r\to 0$ (Figure 1). The softcore energies were then obtained by analytical integration of the polynomial to ensure consistent energy and force data. Without appropriate precautions, this approach can lead to artificial softening of the minimum of the dimer curve if the switching point is too close to the minimum. Conversely, a switching point that is too far from equilibrium would lead to large energy and force values for small separations, which are more challenging for the model to learn.
@@ II.2.2 Solvation free energies in a non-aqueous solvent
We have shown that MACE-OFF24-SC is capable of highly accurate solvation free energy predictions across a range of small organic molecules. However, whilst these serve as a useful and important benchmark, often the more relevant quantity in drug discovery is the distribution coefficient, which additionally requires the calculation of a solvation free energy in a nonpolar solvent. This probes the ability of the model to describe accurately the intermolecular interactions between the solute and the nonpolar organic solvent, in this case octanol. Although the greater conformational flexibility of the solvent makes convergence of the octanol leg of the thermodynamic cycle slightly more challenging, these calculations were still performed in reasonable computational time, requiring at most 48 hours of walltime (see section II.2.4).
In this section, we evaluate the accuracy of MACE-OFF24-SC predictions on several octanol solvation free energies extracted from the MNSol database [57]. Whilst MNSol does not provide experimental uncertainties, we adopt the best practice developed by the FreeSolv authors and assume an experimental error of 0.6 kcal/mol. With this threshold, all MACE-OFF24-SC predictions fall within the experimental error (Figure 6 and Section S2), indicating that the model is indeed capable of successfully predicting these interactions. Here, we compare to the GAFF2 forcefield equipped with the recently published ABCG2 charge model instead of the normal AM1-BCC, which has been tuned to improve the accuracy of hydration and transfer free energy calculations compared to the standard charge model [56]. Hence, this test represents an extremely stringent comparison for MACE-OFF24-SC, since the empirical forcefield has been directly optimised for the task.
@@
Figure 6: Solvation free energies of a series of small organic molecules in octanol. The shaded region represents a 0.6 kcal/mol deviation.
@@ II.2.3 LogP calculations for drug-like molecules
In order to probe the boundaries of the model’s capabilities, we performed full solvation free energy calculations, in both water and octanol solvents, on a selection of drug-like compounds from the CHEMBL database for which an experimental logP value is available via retention time-based UPLC/MS chromatographic analysis. All compounds satisfy Lipinski’s Rule of 5 criteria [58] and are representative of the chemical complexity and flexibility of molecules typical of a medicinal chemistry campaign. Structures of the compounds are shown in Section S9.
@@
Figure 7: Experimental logP for CHEMBL compounds compared to OpenFF 2.1 and MACE-OFF24-SC. *Experimental value recorded as $>5.4$, so quantitative comparison is not possible.
Figure 7 compares both MACE-OFF24-SC and OpenFF 2.1 with experimental measurements. Here, for the first time we see MACE-OFF24-SC significantly outperforming the classical forcefield, with an RMSE of 0.45 log units, compared to 4.02 for OpenFF 2.1. This result is extremely encouraging, suggesting that the model is capable of transferring nonbonded interactions learned from the training set to highly functionalised, conformationally complex molecules. This result validates our hypothesis that machine learning forcefields, which have learned these interactions from first principles, may show superior transferability to drug-like chemical space, compared to the explicit parametrisation approach used to fit empirical forcefields.
@@ III Conclusion
We have introduced a modified version of the MACE-OFF24 transferable organic machine learned forcefield that enables, theoretically rigorous condensed phase alchemical free energy simulations with first principles based machine learned forcefields. A dual approach of softcore repulsion and $\lambda$-dependent two-body interactions enable numerically well-behaved simulations of weakly interacting atoms, directly analogous to softcore formulations of classical forcefields. Our method enables access to quantities directly relevant to drug discovery at ab initio quality.
@@
As a final stringent test of the accuracy of the model, we computed logP values for a series of compounds from the CHEMBL database. These compounds are drug-like in terms of size, complexity and functionality, are expected to be neutral at physiological pH and have experimental logP values. On this benchmark, we significantly outperform OpenFF 2.1, a state-of-the-art classical forcefield, decreasing the RMSE by an order of magnitude. This performance is highly encouraging for future applications of machine learning forcefields in drug discovery, and represents a significant step towards replacing labour and resource intensive chemical synthesis and experiment with fast and accurate molecular simulation.
@@ Acknowledgements
Molecule Experiment Experiment Error MACE MACE Error GAFF2/ABCG2
methanol -3.87 0.6 -3.87 0.09 -3.223
phthalimide -11.18 0.6 -10.25 0.21 -10.945
methane 0.51 0.6 0.26 0.15 0.64
thiophene -3.89 0.6 -3.05 0.11 -3.807
benzonitrile -6.09 0.6 -6.56 0.15 -6.169
caffeine -12.54 0.6 -11.73 0.17 -15.708
bromomethane -2.43 0.6 -1.58 0.2 -2.264
o-cresol -8.49 0.6 -8.95 0.16 -8.58
benzene -3.72 0.6 -3.7 0.11 -3.121
acetone -3.15 0.6 -3.26 0.19 -2.935
Table S3: Solvation free energies in octanol for a series of organic solutes. GAFF2-ABCG2 values are taken from the literature. [56]
@@
Compound Experimental MACE GAFF
mobley_2789243 -4.01 -2.9 -2.77
mobley_1244778 -5.48 -6.44 -4.34
mobley_6091882 1.28 0.33 2.33
mobley_4291494 -4.39 -3.77 -3.87
mobley_8966374 -5.51 -4.65 -6.16
mobley_1881249 -3.52 -4.64 -3.17
mobley_4884177 -4.59 -5.78 -2.96
mobley_6102880 -2.87 -3.93 -2.81
mobley_7039935 -7.67 -7.2 -7.14
mobley_9029594 -0.22 -0.9 0.88
mobley_2782339 -9.61 -8.49 -11.82
mobley_7326982 -4.74 -5.44 -3.43
mobley_8764620 -6.75 -7.82 -5.58
mobley_6353617 -8.26 -7.52 -7.84
mobley_8899867 -5.49 -5.76 -4.29
mobley_6303022 -4.73 -5.62 -3.19
mobley_6794076 -7.62 -7.3 -7.57
mobley_4188615 -9.29 -8.49 -8.11
mobley_1963873 -10.0 -10.43 -8.28
mobley_4639255 -9.3 -10.63 -7.27
mobley_8124669 -8.21 -7.27 -6.85
mobley_7608462 -1.12 -0.48 -0.47
mobley_8578590 -9.28 -10.88 -8.24
mobley_6988468 -4.33 -5.16 -3.31
mobley_6861308 -9.13 -10.39 -11.81
mobley_5220185 -5.45 -5.14 -5.81
mobley_1352110 -3.64 -3.81 -4.57
mobley_2316618 -14.21 -14.34 -13.6
mobley_7010316 -5.46 -6.35 -4.18
mobley_6739648 -5.06 -5.94 -4.27
mobley_1903702 -4.62 -5.22 -3.15
mobley_2049967 -1.5 -2.04 0.39
mobley_9055303 2.0 2.15 2.45
mobley_1636752 -5.1 -5.09 -3.49
mobley_2008055 1.83 1.92 2.46
mobley_2310185 -5.0 -4.98 -3.39
mobley_20524 -6.6 -6.36 -5.71
Table S4: Results for MACE and GAFF on the freesolv subset compared to experiment

### v5

# Computing solvation free energies of small molecules with experimental accuracy

@@ I Introduction
$\Delta G=\int_{0}^{1}\left\langle\frac{\partial H(\vec{r},\lambda)}{\partial\lambda}\right\rangle_{\lambda}d\lambda$ (2)
@@
$U(\lambda,r)=4\epsilon\lambda^{n}\Bigg[\Bigg.\left(\alpha_{LJ}(1-\lambda)^{m}+\left(\frac{r}{\sigma}\right)^{6}\right)^{-2}-\\ \left(\alpha_{LJ}(1-\lambda)^{m}+\left(\frac{r}{\sigma}\right)^{6}\right)^{-1}\Bigg$$\Bigg$] (3)
@@
Although MLPs can deliver significant accuracy improvements compared to empirical forcefields, their computational expense and the high sampling requirements of free energy calculations have meant that their application has been mostly restricted to corrective perturbations [17, 55, 41]. In this approach, the intramolecular interactions of a subset of atoms (usually the organic molecule) are modelled by the MLP, while the remaining bonded interactions and all nonbonded interactions are modelled by the empirical forcefield. Although there is some evidence that this approach can modestly improve the accuracy of RBFE calculations, it does not address the parametrisation of nonbonded interactions, which are known to be crucial for obtaining accurate free energies [55]. Indeed, it has recently been shown that, for hydration free energies, MLP corrections with mechanical embedding fail to provide statistically significant accuracy improvements [41]. It is worth noting that more sophisticated electrostatic embedding ML/MM schemes have recently been proposed [68, 51, 54]. However, the performance of these methods in alchemical binding free energy calculations has not yet been established.
@@ II Results & Discussion
II.1 Alchemical simulations with MLPs
@@ II.1.1 Softcore dimer curves
[…]
@@ II.2.2 Solvation free energies in a non-aqueous solvent
We have shown that MACE-OFF24-SC is capable of highly accurate solvation free energy predictions across a range of small organic molecules. However, whilst these serve as a useful and important benchmark, often the more relevant quantity in drug discovery is the distribution coefficient, which additionally requires the calculation of a solvation free energy in a nonpolar solvent. This probes the ability of the model to describe accurately the intermolecular interactions between the solute and the nonpolar organic solvent, in this case octanol. Although the greater conformational flexibility of the solvent makes convergence of the octanol leg of the thermodynamic cycle slightly more challenging, these calculations were still performed in reasonable computational time, requiring at most 48 hours of walltime (see section II.2.4). We first confirmed that MACE-OFF24-SC is capable of capturing the basic thermodynamic properties of liquid octanol, resulting a density within 5% of the experimental value (Figure S11).
In this section, we evaluate the accuracy of MACE-OFF24-SC predictions on 10 octanol solvation free energies extracted from the MNSol database [47]. Whilst MNSol does not provide experimental uncertainties, we adopt the best practice developed by the FreeSolv authors and assume an experimental error of 0.6 kcal/mol. With this threshold, all MACE-OFF24-SC predictions fall within the experimental error (Figure 6 and Section S2), indicating that the model is indeed capable of successfully predicting these interactions. Here, we compare to the GAFF2 forcefield equipped with the recently published ABCG2 charge model instead of the normal AM1-BCC, which has been tuned to improve the accuracy of hydration and transfer free energy calculations compared to the standard charge model [37]. Hence, this test represents an extremely stringent comparison for MACE-OFF24-SC, since the empirical forcefield has been directly optimised for the task.
@@
Figure 6: Solvation free energies of a series of 10 small organic molecules in octanol. The shaded region represents a 0.6 kcal/mol deviation.
@@ II.2.3 LogP calculations for drug-like molecules
In order to probe the boundaries of the model’s capabilities, we performed full solvation free energy calculations, in both water and octanol solvents, on a selection of 16 drug-like compounds from the CHEMBL database for which an experimental logP value is available via retention time-based UPLC/MS chromatographic analysis. All compounds satisfy Lipinski’s Rule of 5 criteria [44] and are representative of the chemical complexity and flexibility of molecules typical of a medicinal chemistry campaign. Structures of the compounds are shown in Section S9.
@@
Figure 7: Experimental logP for CHEMBL compounds compared to OpenFF 2.1, MACE-OFF24-SC and GAFF2/ABCG2. *Experimental value recorded as >5.4>5.4, so quantitative comparison is not possible.
Figure 7 compares both MACE-OFF24-SC and OpenFF 2.1 with experimental measurements. Here, for the first time we see MACE-OFF24-SC significantly outperforming the classical forcefields, with an RMSE of 0.45 log units, compared to 4.02 for OpenFF 2.1 and 8.64 for GAFF2/ABCG2. This result is extremely encouraging, suggesting that the model is capable of transferring nonbonded interactions learned from the training set to highly functionalised, conformationally complex molecules. This result validates our hypothesis that machine learning forcefields, which have learned these interactions from first principles, may show superior transferability to drug-like chemical space, compared to the explicit parametrisation approach used to fit empirical forcefields [34]. Given the accurate predictions of GAFF2/ABCG2 on small molecule octanol solvation free energies, this suggests that the error arises in the transferability of the parameters from small molecules on which the nonbonded parameters are fitted to larger, more complex and conformationally flexible systems. These results are in qualitative agreement with other recently published benchmarking studies, which show that, whilst ABCG2 charges lead to improved performance on hydration free energies of small molecules, they fail to outperform AM1-BCC charges on closely related protein-ligand binding free energy calculations, further highlighting that parameter transferability remains a significant challenge for classical small molecule forcefields [7].
@@ III Conclusion
We have introduced a modified version of the MACE-OFF24 transferable organic MLPs that enables, theoretically rigorous condensed phase alchemical free energy simulations with first principles based MLPs. A dual approach of softcore repulsion and $\lambda$-dependent two-body interactions enable numerically well-behaved simulations of weakly interacting atoms, directly analogous to softcore formulations of classical forcefields. Our method enables access to quantities directly relevant to drug discovery at ab initio quality.
@@
As a final stringent test of the accuracy of the model, we computed logP values for a series of compounds from the CHEMBL database. These compounds are drug-like in terms of size, complexity and functionality, are expected to be neutral at physiological pH and have experimental logP values. On this benchmark, we report an order of magnitude improvement in RMSE compared to OpenFF 2.1, a state-of-the-art classical forcefield, and GAFF2/ABCG2, which utilises a recently developed charge partitioning scheme [37]. This performance is highly encouraging for future applications of machine learning forcefields in drug discovery, and represents a significant step towards replacing labour and resource intensive chemical synthesis and experiment with fast and accurate molecular simulation.
@@ IV.3 Empirical forcefield solvation free energies
GAFF2/ABCG2 calculations were performed with GROMACS 2024.1. Parameters were assigned with Antechamber. Systems were energy minimised for a maximum of 50000 steps and subsequently equilibrated for 1 ns in the NVT ensemble and the NPT ensemble for 1 ns. 8 and 16 windows were used to annihilate the Coulomb and Van der Waals interactions respectively. Each alchemical window was run for 500 ps. An example MDP file for the production free energy calculations is included in the section S11. Free energy estimates were calculated using gmx bar.
@@ Acknowledgements
Data Availability
An example script to reproduce the replica exchange calculations is provided in the SI. The mace-md package used to run the simulations is available at https://github.com/jharrymoore/mace-md/tree/master. Calculated solvation free energy values for all forcefields are available in the supplementary information.
@@
Molecule SMILES Experiment Experiment Error MACE MACE Error GAFF2/ABCG2
methanol CO -3.87 0.6 -3.87 0.09 -3.223
phthalimide O=C1NC(=O)c2ccccc12 -11.18 0.6 -10.25 0.21 -10.945
methane C 0.51 0.6 0.26 0.15 0.64
thiophene c1ccsc1 -3.89 0.6 -3.05 0.11 -3.807
benzonitrile N#Cc1ccccc1 -6.09 0.6 -6.56 0.15 -6.169
caffeine Cn1cnc2c1c(=O)n(C)c(=O)n2C -12.54 0.6 -11.73 0.17 -15.708
bromomethane CBr -2.43 0.6 -1.58 0.2 -2.264
o-cresol Cc1ccccc1O -8.49 0.6 -8.95 0.16 -8.58
benzene c1ccccc1 -3.72 0.6 -3.7 0.11 -3.121
acetone CC(=O)C -3.15 0.6 -3.26 0.19 -2.935
Table S3: Solvation free energies in kcal/mol in octanol for a series of organic solutes. GAFF2-ABCG2 values are taken from the literature. [37]
@@
Compound Name Exp MACE GAFF OpenFF Exp Error MACE Error GAFF Error OpenFF Error
cycloheptanol -5.48 -6.44 -4.34 -4.59 0.60 0.18 0.03 0.19
ethylene 1.28 0.33 2.33 2.29 0.60 0.09 0.01 0.09
2-chloropyridine -4.39 -3.77 -3.87 -3.81 0.60 0.14 0.03 0.16
2-methylpyrazine -5.51 -4.65 -6.16 -6.55 0.60 0.15 0.03 0.18
pentan-2-one -3.52 -4.64 -3.17 -3.15 0.60 0.15 0.03 0.20
3-ethylpyridine -4.59 -5.78 -2.96 -3.28 0.60 0.24 0.03 0.17
diphenyl ether -2.87 -3.93 -2.81 -3.23 0.69 0.27 0.03 0.19
naphthalen-1-ol -7.67 -7.20 -7.14 -8.43 0.60 0.19 0.03 0.20
fluoromethane -0.22 -0.90 0.88 0.93 0.60 0.13 0.01 0.10
phthalimide -9.61 -8.49 -11.82 -11.73 0.50 0.15 0.03 0.22
propan-2-ol -4.74 -5.44 -3.43 -4.00 0.60 0.14 0.02 0.16
pyridine-3-carbonitrile -6.75 -7.82 -5.58 -5.93 0.60 0.17 0.03 0.17
1-(3-pyridyl)ethanone -8.26 -7.52 -7.84 -7.09 0.60 0.20 0.03 0.20
cyclopentanol -5.49 -5.76 -4.29 -4.32 0.60 0.15 0.03 0.17
4-ethylpyridine -4.73 -5.62 -3.19 -4.29 0.60 0.15 0.03 0.17
1-(4-pyridyl)ethanone -7.62 -7.30 -7.57 -7.42 0.60 0.32 0.03 0.21
methane 2.00 1.33 2.45 2.40 0.20 0.09 0.01 0.15
ethanol -5.00 -5.83 -3.39 -4.98 0.60 0.16 0.02 0.12
phenol -6.60 -6.29 -5.71 -6.36 0.20 0.13 0.03 0.15
ethane 1.83 1.02 2.46 1.92 0.60 0.12 0.01 0.14
methanol -5.10 -5.61 -3.49 -5.09 0.60 0.16 0.02 0.17
N,N-dimethylbenzamide -9.29 -8.49 -8.11 -7.93 0.60 0.27 0.03 0.21
N-methylacetamide -10.00 -10.43 -8.28 -7.90 0.60 0.31 0.02 0.18
ethylene glycol -9.30 -10.63 -7.27 -8.89 0.60 0.15 0.03 0.20
alachlor -8.21 -7.27 -6.85 -8.28 0.29 0.26 0.05 0.25
chlorobenzene -1.12 -0.48 -0.47 -1.06 0.60 0.15 0.02 0.15
methylsulfinylmethane -9.28 -10.88 -8.24 -7.03 0.57 0.20 0.02 0.17
2-ethylpyridine -4.33 -5.16 -3.31 -3.37 0.60 0.16 0.03 0.19
fenuron -9.13 -10.39 -11.81 -10.81 1.93 0.17 0.04 0.23
2-ethylpyrazine -5.45 -5.14 -5.81 -5.94 0.60 0.19 0.03 0.21
pebulate -3.64 -3.81 -4.57 -4.13 1.93 0.24 0.04 0.23
1-(2-hydroxyethylamino)- 9,10-anthraquinone -14.21 -14.34 -13.60 -14.45 1.10 0.34 0.05 0.28
cyclohexanol -5.46 -6.35 -4.18 -3.83 0.60 0.14 0.03 0.21
1,4-dioxane -5.06 -5.94 -4.27 -3.58 0.60 0.16 0.02 0.16
butan-2-ol -4.62 -5.22 -3.15 -3.37 0.60 0.18 0.03 0.17
methylsulfanylethane -1.5 -2.04 0.39 0.49 0.2 0.15 0.02 0.14
Table S4: Solvation free energies in kcal/mol for MACE, OpenFF and GAFF on the FreeSolv subset compared to experiment.
@@
Listing 2: Script to run equilibrium hydration free energy calculation with openff-2.1.0.
@@
Appendix S10 Octanol Density
Figure S11: Density of liquid octanol simulated by MACE-OFF24-SC. Average density was calculated from the last 20 ps of the simulation
Appendix S11 Example GROMACS MDP file
1
2title = Absolute Solvation Free Energy in Octanol
3integrator = sd
4nsteps = 250000 ; 500 ps per window
5dt = 0.002
6nstenergy = 500
7nstlog = 500
8nstxout-compressed = 500
9
10continuation = yes
11constraint_algorithm = lincs
12constraints = h-bonds
13lincs_iter = 1
14lincs_order = 4
15
16nstlist = 40
17cutoff-scheme = Verlet
18ns_type = grid
19coulombtype = PME
20rcoulomb = 1.2
21rvdw = 1.2
22pbc = xyz
23disp_corr = EnerPres
24
25tc-grps = System
26tau_t = 0.1
27ref_t = 298.15
28
29pcoupl = Parrinello-Rahman
30pcoupltype = isotropic
31tau_p = 2.0
32ref_p = 1.0
33compressibility = 4.5e-5
34
35; Free energy control
36free_energy = yes
37init_lambda_state = 0 ; will be overwritten
38delta_lambda = 0
39calc_lambda_neighbors = 1 ; calculate delta H to neighbors
40
41; Lambda vectors
42coul_lambdas = 0.00 0.14 0.29 0.43 0.57 0.71 0.86 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
43vdw_lambdas = 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.06 0.12 0.19 0.25 0.31 0.37 0.44 0.50 0.56 0.62 0.69 0.75 0.81 0.87 0.94 1.00
44
45couple-moltype = mol_009
46couple-lambda0 = vdw-q
47couple-lambda1 = none
48couple-intramol = yes
49
50sc-alpha = 0.5
51sc-power = 1
52sc-sigma = 0.3
53nstdhdl = 100
Listing 3: mdp file for production GROMACS calculations for GAFF2/ABCG2 calculations